Find all the units of Z[ √− 7].


1
Expert's answer
2021-05-07T09:18:14-0400

Take x, y ∈ Z[ √− 7] such that x has the form of a+ b √− 7 while y takes the form c + d √− 7 for the arbitrary values of x and y.

if (a+ b √− 7) (c + d √− 7) = 1 then (a - b √− 7) (c - d √− 7) =1.

Therefore;

(a+ b √− 7) (c + d √− 7)(a - b √− 7) (c - d √− 7) = 1 since 1*1 = 1

thus;

(a2 + 7b2)(c2 + 7d2) = 1 hence

b2 = d2 = 0 and

a2 = c2 = 1 or -1 thus a=c = 1 or -1 which are the units in Z[ √− 7]